You buy a computer directly from the manufacturer for R73320.00 and Agree to repay it in equal installments over three at end of each month,starting one month from now. The interest rate is 10,7% per year, compounded monthly. How much interest will you pay in total?


1
Expert's answer
2021-05-12T05:08:53-0400

Given the following:

FV= R7332

Rate(r)= 10.7% or 0.107

Time(t)=3years

The installmental payment for each year is given as:

"PV=\frac{FV}{[\frac{(1+r)^t-1}{r}]}"


"PV=\frac{7332}{[\frac{(1+r)^3-1}{0.107}]}"


"PV=R2200.18"





The above picture shows the payment schedule for the repayment of the loan.


Year 1=0

Year 2 "=0.107\u00d7R2200.18"

"=R235.42"

Year 3 "=0.107\u00d7R4635.7"

"=R496.02"

Total interest to be repaid on the loan of R7332 is R731.44

= "0+R235.42+R496.02=R731.44"